# The galaxy-IGM connection in thesan: observability and information content of the galaxy-Lyman- cross-correlation at z6

### v1

@@ Abstract
The galaxyāLyman- cross-correlation (GaLCC) is a promising tool to study the interplay of galaxies and inter-galactic medium (IGM) in the first billion years of the Universe. Here we thoroughly characterise the impact of observational limitations on our ability to retrieve the intrinsic GaLCC and provide new physical insights on its origin and connection to other IGM properties. This is extremely relevant to identify promising datasets, design future surveys and assess the limitations of current measurements. We find that sightline-to-sightline variations demand at least 25 independent sightlines to quantitatively recover the true signal. Once this condition is met, the intrinsic signal can be recovered even for a relatively low signal-to-noise ratio and spectral resolution. The galaxy selection method does not affect the inferred GaLCC and lightcone effects are relevant whenever observations span a redshift window broader than . We discuss the implication for previous theoretical studies that did not account for them. We elucidate explicitly for the first time the physical origin of the GaLCC and demonstrate that this signal is collectively sourced by the ensemble of galaxies residing in overdense regions rather than individual objects. We show that the GaLCC measured for opaque sightlines shows a larger peak at smaller scales with respect to transparent lines of sight. We connect this to the evolution of the mean free path of ionizing photons, showing that the GaLCC peak position has a very similar evolution but on smaller scales, as it probes only the core of ionised regions. Finally, we discuss which ongoing surveys can be used to measure the GaLCC and provide an initial analysis of future developments, including using galaxies as background sources, and the application to helium reionization. Our results outline a bright future for the GaLCC as a tool to unveil the galaxy-IGM interplay during the first billion years of the Universe.
@@ 1 Introduction
Following the emergence of the first stars and galaxies few hundred millions years after the Big Bang, the ultraviolet photons emitted by such objects began to ionize the intergalactic medium (IGM) gas between them in what is known as the Epoch of cosmic Reionization (EoR). These ionised regions eventually grew to encompass the entire Universe, marking the end of such epoch.
@@
The study of cosmic reionization is underpinned by a growing body of observational constraints, that just in the last few years started to probe beyond the tail end of this process (for a somewhat up-to-date collection see e.g.Ā corecon.readthedocs.io/, Garaldi, 2023). These include global constraints from the cosmic microwave background (CMB, e.g.Ā Planck Collaboration etĀ al., 2020; Pagano etĀ al., 2020; de Belsunce etĀ al., 2021) and localised (in time and space) constraints through āĀ among othersĀ ā the Lyman- (Ly) absorption in quasar spectra (e.g.Ā Fan etĀ al., 2006; McGreer etĀ al., 2011; Yang etĀ al., 2020; Lu etĀ al., 2020; Bosman etĀ al., 2022), the visibility evolution of galaxies (e.g.Ā Ota etĀ al., 2008; Pentericci etĀ al., 2014; Mesinger etĀ al., 2014), as well as the damping wing of both quasars (e.g.Ā Mortlock etĀ al., 2011; Greig etĀ al., 2017; Wang etĀ al., 2020; ÄurovÄĆkovĆ” etĀ al., 2024) and Gunn-Peterson troughs (Spina etĀ al., 2024; Zhu etĀ al., 2024).
@@
The launch of the James Webb Space Telescope (JWST) has enabled a fast and dramatic progress on the characterisation of the sources of reionization. Thanks to its near-infrared capabilities and sensitivity, it has already unveiled the properties of a large number of ānormalā galaxies within the first billion years of the Universe (e.g.Ā Finkelstein etĀ al., 2023; Matthee etĀ al., 2023; Harikane etĀ al., 2023; Eisenstein etĀ al., 2023; Meyer etĀ al., 2024). This is not only extremely important to progress further in our understanding of primeval galaxy formation and cosmic reionization, but allows us for the first time to observationally study how these two processes influence each other in the reionizaing Universe.
@@
In recent years, the cross-correlation between the transmitted flux in the Lyman- (Ly) forest and the position of galaxies around the line of sight (hereafter named galaxy-Ly cross-correlation, or GaLCC) has been used to probe the complex galaxy-IGM interplay during the EoR (starting with Kakiichi etĀ al., 2018). This quantity shows two prominent features, namely an excess of transmitted Ly flux at distances from galaxies and a strong suppression of such flux at (Meyer etĀ al., 2019, 2020). The former has been interpreted as a (transverse) proximity effect driven by the ionizing radiation field of the galaxies, while the latter is typically ascribed to the overdensity in which galaxies reside boosting the hydrogen recombination rate. The position and amplitude of the flux excess strongly depends on the progress of reionization (Garaldi etĀ al., 2022), rendering it a powerful tool to constrain the timing of reionization. Little is known beyond this, since theoretical studies of the GaLCC are very limited, mainly as a consequence of the simultaneous requirements of scales to properly capture the inhomogeneous reionization process (Iliev etĀ al., 2014; Kaur etĀ al., 2020; Gnedin & Madau, 2022) while resolving galactic properties (since a failure to do so can erase the GaLCC, Garaldi etĀ al., 2019, 2022).
@@
Observationally studying the GaLCC is also difficult due to the simultaneous requirements of spectroscopy and galaxy detection around the line of sight. This has so far limited studies of this cross-correlation to a small number of sightlines. The status quo is however rapidly changing thanks to the JWST. Recently, the EIGER program published their measurements of the GaLCC in Kashino etĀ al. (2023), based on the first observed line of sight (out of six planned). Intriguingly, the comparison of their results with numerical predictions (from Garaldi etĀ al., 2022) seems to indicate that the simulated a very late end of the EoR. The ASPIRE (Wang etĀ al., 2023) program is also measuring the GaLCC (K. Kakiichi, private communication), thanks to its NIRCam/WFSS observations of 25 quasar fields in the redshift range . Despite such progress, available observations remain sparse and with significant differences in their galaxy identification method (e.g.Ā CāIV absorption, OāIII emission, Lyman-break, etc.), quasar spectrum noise level and resolution, and field coverage. However, the impact of these differences has not yet been investigated, nor has been determined a set of minimal requirements for observations of the GaLCC.
@@
In this paper we provide a thorough characterization of the impact on the inferred GaLCC of a broad range of observational limitations and choices. This enables us to assess the reliability and guide the interpretation of current and future studies of this quantity, as well as to determine a set of minimal requirements for future observations of the GaLCC. We describe the simulation set used and the production of synthetic observations in Sec.Ā 2. In Sec.Ā 3 we present our results concerning the impact of observational limitations on the inferred GaLCC, while in Sec.Ā 4 we connect the properties of the (simulated) signal to other physical properties of the Universe. Finally, in Sec.Ā 5 we discuss potential future developments of the GaLCC, including the prospect of employing ongoing surveys to extend observations of this quantity, and we present concluding remarks in Sec.Ā 6. All quantities are in comoving units throughout the paper, unless specified otherwise.
@@ 2 Methods
Faithfully studying the interplay between galaxies and the large-scale reionization of the inter-galactic medium is challenging, as it requires simultaneous modeling of sub-galactic and inter-galactic scales in volumes at least (Gnedin & Madau, 2022, but potentially much larger, see Iliev etĀ al. 2014), accounting for the inhomogeneous reionization process and including realistic galaxy formation physics. This is a formidable challenge, that has been conquered only by a handful of simulations to date (Ocvirk etĀ al., 2016, 2020; Lewis etĀ al., 2022; Aubert etĀ al., 2018; Gnedin, 2014). In this work we employ one of them, the thesan suite (Kannan etĀ al., 2022a; Garaldi etĀ al., 2022; Smith etĀ al., 2022a; Garaldi etĀ al., 2024), which we briefly describe in the following and refer the interested reader to Kannan etĀ al. (2022a) and Garaldi etĀ al. (2024) for a more thorough description of the numerical and physical setup, as well as of the numerical products available.
@@ 2.1 The thesan simulations
The thesan simulation suite is a set of radiation-hydrodynamical simulations, recently made publicly available at www.thesan-project.com (Garaldi etĀ al., 2024). They are designed following two main tenets, i.e.Ā (i) to simultaneously capture IGM and galactic properties during the EoR while (ii) minimizing the number of free parameters at . In order to do so, the thesan simulations employ the successful IllustrisTNG galaxy formation model (Weinberger etĀ al., 2017; Pillepich etĀ al., 2018), coupled to the dust model of McKinnon etĀ al. (2016) and the arepo-rt (Kannan etĀ al., 2019) radiation transport module of the arepo code (Springel, 2010; Weinberger etĀ al., 2020). By maintaining the free parameters of these models fixed to the values calibrated on the low- Universe, thesan has a single (additional) free parameter, namely the escape fraction of ionizing photons from the birth cloud of the star.111This is not the escape fraction typically discussed in the context of cosmic reionization, which refers to the ionizing photons escape from the entire galaxy/halo and that the simulations can predict (Yeh etĀ al., 2023). Rather, this represents the absorption of ionizing photons by unresolved structures around the birth place of stars, typically on scales of few tens of parsecs (see e.g.Ā Table 1 of Garaldi etĀ al., 2024). The latter is calibrated by requiring the simulations to approximately match the observed ālateā reionization history (e.g.Ā Zhu etĀ al., 2020; Bosman etĀ al., 2022; Kulkarni etĀ al., 2019; Keating etĀ al., 2020; Nasir & DāAloisio, 2020).
@@
All thesan simulations employ a Planck Collaboration etĀ al. (2016) cosmology and have a box size of . The flagship simulation, thesan-1, has a mass resolution sufficient to resolve atomic cooling haloes, the smallest of structures significantly contributing to the ionizing photon budget (for a recent confirmation of the negligible role of mini-haloes see e.g.Ā Gnedin, 2024). This is the simulation used in this paper. thesan matches the main galaxy properties observed in the pre-JWST (Kannan etĀ al., 2022a) and JWST (Garaldi etĀ al., 2024; Shen etĀ al., 2024) era, as well the observed IGM properties (Garaldi etĀ al., 2022).
@@
It should be noted that, despite being one of the largest-volume radiation-hydrodynamical simulations of the Universe currently able to capture galaxy properties (only comparable to CROC āĀ Gnedin 2014Ā ā and CoDaIII āĀ Lewis etĀ al. 2022), the volume covered remains somewhat small, and only barely approaching the volumes needed for a converged reionization history (e.g.Ā Iliev etĀ al., 2014; Kaur etĀ al., 2020; Gnedin & Madau, 2022). For instance, Becker etĀ al. (2011) reported the discovery of a very long and very opaque Gunn-Peterson trough in an otherwise mostly ionised Universe. Extreme features like this one can not be captured in our simulations (see e.g.Ā Keating etĀ al. 2020 for an estimation of the volume needed to simulate similar features). Therefore, it is possible that our results will be marginally affected by a somewhat suppressed variability due to the inability of thesan to capture the most extreme features. However, we expect these to be very rare, and therefore to bear a small impact on our results.
@@ 2.2 Synthetic spectra
The main data product used in this paper are synthetic lines of sights (LOS) extracted from the simulation outputs (see Section 3.10 of Garaldi etĀ al., 2024). These are extracted using the colt code (last described in Smith etĀ al., 2022b), which uses the native Voronoi tessellation of the simulation to retain the full spatial information available. Using the gas properties extracted in this way, we construct synthetic Ly forest spectra using the full Voigt-Hjerting line profile (Hjerting 1938, through the approximation of Harris 1948 and Tepper-GarcĆa 2006). We include the effects of gas temperature and peculiar velocities. Our spectra have a spectral resolution of . For each simulation snapshot (i.e.Ā approximately every 11 Myr) having redshift between , we produce 300 lines of sight (in addition to the 150 already available in the public data release). These new LOS originate from a random location in the plane, are aligned along the direction and are long. We note here that, thanks to the unstructured mesh of arepo, the synthetic spectra do not suffer from any grid-alignment artefact.
For this paper, we combine the LOS produced at each simulation snapshot to produce lightcone spectra, employing a piecewise-constant approximation. We release these sightlines on the thesan website (see AppendixĀ A for more information). In Fig.Ā 1 we show an example of such lightcone LOS. The background map shows the density distribution within 30 Mpc of the sightline (along the direction of the simulation, left-hand side vertical axis and top left colorbar). The saturation of this map reflects the amount of neutral hydrogen in the IGM (with grey regions corresponding to and progressively more saturated colors indicating lower neutral fractions, top right colorbar). The white spectrum super-imposed to this map shows the normalised transmitted flux along the LOS (right-hand side y axis). The evolving redshift of the spectrum is indicated by the horizontal axis. This visually shows the well-known boost in Ly transmission in highly-ionised gas, which is preferentially found at low redshift.
In the paper, when not stated otherwise, the cross-correlation is always computed using all the 300 sightlines (either at fixed redshift or using the lightcone LOS) and considering galaxies with stellar mass . Additionally, all quantities are expressed in comoving units, unless specified otherwise.

### v2

# The galaxy–IGM connection in thesan: observability and information content of the galaxy–Lyman- cross-correlation at z6

@@ Abstract
The galaxy–Lyman- cross-correlation (GaLCC) is a promising tool to study the interplay of galaxies and inter-galactic medium (IGM) in the first billion years of the Universe. Here we thoroughly characterise the impact of observational limitations on our ability to retrieve the intrinsic GaLCC and provide new physical insights on its origin and connection to other IGM properties. This is extremely relevant to identify promising datasets, design future surveys and assess the limitations of current measurements. We find that sightline-to-sightline variations demand at least 25 independent sightlines to quantitatively recover the true signal. Once this condition is met, the intrinsic signal can be recovered even for a relatively low signal-to-noise ratio and spectral resolution. The galaxy selection method does not affect the inferred GaLCC and lightcone effects are relevant whenever observations span a redshift window broader than . We discuss the implication for previous theoretical studies that did not account for them. We elucidate explicitly for the first time the physical origin of the GaLCC and demonstrate that this signal is collectively sourced by the ensemble of galaxies residing in overdense regions rather than individual objects. We show that the GaLCC measured for opaque sightlines shows a larger peak at smaller scales with respect to transparent lines of sight. We connect this to the evolution of the mean free path of ionizing photons, showing that the GaLCC peak position has a very similar evolution but on smaller scales, as it probes only the core of ionized regions. Finally, we discuss which ongoing surveys can be used to measure the GaLCC and provide an initial analysis of future developments, including using galaxies as background sources, and the application to helium reionization. Our results outline a bright future for the GaLCC as a tool to unveil the galaxy–IGM interplay during the first billion years of the Universe.
@@ 1 Introduction
Following the emergence of the first stars and galaxies within a few hundred million years after the Big Bang, the ultraviolet photons emitted by such objects began to ionize the intergalactic medium (IGM) gas between them in what is known as the Epoch of cosmic Reionization (EoR). These ionized regions eventually grew to encompass the entire Universe, marking the end of such epoch.
@@
The study of cosmic reionization is underpinned by a growing body of observational constraints, that just in the last few years started to probe beyond the tail end of this process (for a community-updated collection see e.g. corecon.readthedocs.io; Garaldi, 2023). These include global constraints from the cosmic microwave background (CMB, e.g. Planck Collaboration et al., 2020; Pagano et al., 2020; de Belsunce et al., 2021) and localised (in time and space) constraints through – among others – the Lyman- (Ly) absorption in quasar spectra (e.g. Fan et al., 2006; McGreer et al., 2011; Yang et al., 2020; Lu et al., 2020; Bosman et al., 2022), the visibility evolution of galaxies (e.g. Ota et al., 2008; Pentericci et al., 2014; Mesinger et al., 2014), as well as the damping wing of both quasars (e.g. Mortlock et al., 2011; Greig et al., 2017; Wang et al., 2020; Ďurovčíková et al., 2024) and Gunn-Peterson troughs (Spina et al., 2024; Zhu et al., 2024).
@@
The launch of the James Webb Space Telescope (JWST) has enabled a fast and dramatic progress on the characterisation of the sources of reionization. Thanks to its near-infrared capabilities and sensitivity, it has already unveiled the properties of a large number of ‘normal’ galaxies within the first billion years of the Universe (e.g. Finkelstein et al., 2023; Matthee et al., 2023; Harikane et al., 2023; Eisenstein et al., 2023; Meyer et al., 2024). This is not only extremely important to progress further in our understanding of primeval galaxy formation and cosmic reionization, but allows us for the first time to observationally study how these two processes influence each other in the reionizaing Universe.
@@
In recent years, the cross-correlation between the transmitted flux in the Lyman- (Ly) forest and the position of galaxies around the line of sight (hereafter named galaxy–Ly cross-correlation, or GaLCC) has been used to probe the complex galaxy–IGM interplay during the EoR (starting with Kakiichi et al., 2018). This quantity shows two prominent features, namely an excess of transmitted Ly flux at distances from galaxies and a strong suppression of such flux at (Meyer et al., 2019, 2020). The former has been interpreted as a (transverse) proximity effect driven by the ionizing radiation field of the galaxies, while the latter is typically ascribed to the overdensity in which galaxies reside boosting the hydrogen recombination rate. The position and amplitude of the flux excess strongly depends on the progress of reionization (Garaldi et al., 2022), rendering it a powerful tool to constrain the timing of reionization. Little is known beyond this, since theoretical studies of the GaLCC are very limited, mainly as a consequence of the simultaneous requirements of scales to properly capture the inhomogeneous reionization process (Iliev et al., 2014; Kaur et al., 2020; Gnedin & Madau, 2022) while resolving galactic properties (since a failure to do so can erase the GaLCC, Garaldi et al., 2019, 2022). During the revision process of this manuscript, Conaboy et al. (2025) published a thorough investigation of the GaLCC using the Sherwood-relics simulations (Puchwein et al., 2023), which employ a two-step process to include the impact of gas photo-ionisation and photo-heating in the simulation without performing on-the-fly radiation transport.
@@
Observationally studying the GaLCC is also difficult due to the simultaneous requirements of spectroscopy and galaxy detection around the line of sight. This has so far limited studies of this cross-correlation to a small number of sightlines. The status quo is however rapidly changing thanks to the JWST. Recently, the EIGER program published their measurements of the GaLCC in Kashino et al. (2023), based on the first observed line of sight (out of six planned). Intriguingly, matching their results to numerical predictions (from Garaldi et al., 2022) requires a very late end of the EoR. Interestingly, this is in contrast to what is inferred by matching the observations of Meyer et al. (2019) to the simulations in Conaboy et al. (2025). The ASPIRE (Wang et al., 2023) program is also measuring the GaLCC, thanks to its NIRCam/WFSS observations of 25 quasar fields in the redshift range . Recently, results from the first 5 quasar fields were published (Kakiichi et al., 2025), demonstrating an overall good agreement with the thesan simulations. Nevertheless, a detailed comparison revealed some differences pointing towards larger ionised bubbles and stronger UV background and temperature fluctuations. Despite such progress, available observations remain sparse and with significant differences in their galaxy identification method (e.g. C IV absorption, O III emission, Lyman-break, etc.), quasar spectrum noise level and resolution, and field coverage. However, the impact of these differences has not yet been investigated, nor has been determined a set of minimal requirements for observations of the GaLCC.
@@
In this paper we provide a thorough characterization of the impact on the inferred GaLCC of a broad range of observational limitations and choices. This enables us to assess the reliability and guide the interpretation of current and future studies of this quantity, as well as to determine a set of minimal requirements for future observations of the GaLCC. We describe the simulation set used and the production of synthetic observations in Sec. 2. In Sec. 3 we present our results concerning the impact of observational limitations on the inferred GaLCC, while in Sec. 4 we connect the properties of the (simulated) signal to other physical properties of the Universe. Finally, in Sec. 5 we discuss potential future developments of the GaLCC, including the prospect of employing ongoing surveys to extend observations of this quantity, and we present concluding remarks in Sec. 6.
@@ 2 Methods
Faithfully studying the interplay between galaxies and the large-scale reionization of the IGM is challenging, as it requires simultaneous modeling of sub-galactic and inter-galactic scales in volumes at least (Gnedin & Madau, 2022, but potentially much larger, see Iliev et al. 2014), accounting for the inhomogeneous reionization process and including realistic galaxy formation physics. This is a formidable challenge, that has been conquered only by a handful of simulations to date (Ocvirk et al., 2016, 2020; Lewis et al., 2022; Aubert et al., 2018; Gnedin, 2014). In this work we employ one of them, the thesan suite (Kannan et al., 2022a; Garaldi et al., 2022; Smith et al., 2022a; Garaldi et al., 2024), which we briefly describe in the following and refer the interested reader to Kannan et al. (2022a) and Garaldi et al. (2024) for a more thorough description of the numerical and physical setup, as well as of the numerical products available.
@@ 2.1 The thesan simulations
The thesan simulation suite is a set of radiation-hydrodynamical simulations, recently made publicly available at www.thesan-project.com (Garaldi et al., 2024). They are designed following two main tenets, i.e. (i) to simultaneously capture IGM and galactic properties during the EoR while (ii) minimizing the number of free parameters at . In order to do so, the thesan simulations employ the successful IllustrisTNG galaxy formation model (Weinberger et al., 2017; Pillepich et al., 2018), coupled to the dust model of McKinnon et al. (2016) and the arepo-rt (Kannan et al., 2019) radiation transport module of the arepo code (Springel, 2010; Weinberger et al., 2020). By maintaining the free parameters of these models fixed to the values calibrated on the low- Universe, thesan has a single (additional) free parameter, namely the escape fraction of ionizing photons from the birth cloud of the star.111This is not the escape fraction typically discussed in the context of cosmic reionization, which refers to the ionizing photons escape from the entire galaxy/halo and that the simulations can predict (Yeh et al., 2023). Rather, this represents the absorption of ionizing photons by unresolved structures around the birth place of stars, typically on scales of parsecs (see e.g. Table 1 of Garaldi et al., 2024). The latter is calibrated by requiring the simulations to approximately match the observed ‘late’ reionization history (e.g. Zhu et al., 2020; Bosman et al., 2022; Kulkarni et al., 2019; Keating et al., 2020; Nasir & D’Aloisio, 2020).
@@
All thesan simulations employ a Planck Collaboration et al. (2016) cosmology and have a box size of . The flagship simulation, thesan-1, has a mass resolution sufficient to resolve atomic cooling haloes, the smallest of structures significantly contributing to the ionizing photon budget (for a recent confirmation of the negligible role of mini-haloes see e.g. Gnedin, 2024). This is the simulation used in this paper. thesan matches the main galaxy properties observed in the pre-JWST (Kannan et al., 2022a) and JWST (Garaldi et al., 2024; Shen et al., 2024) era, as well the observed IGM properties (Garaldi et al., 2022; Borrow et al., 2023; Neyer et al., 2024; Jamieson et al., 2025; Zhao et al., 2025).
@@
It should be noted that, despite being one of the largest-volume radiation-hydrodynamical simulations of the Universe currently able to capture galaxy properties (only comparable to CROC – Gnedin 2014 – and CoDaIII – Lewis et al. 2022), the volume covered remains somewhat small, and only barely approaching the volumes needed for a converged reionization history (e.g. Iliev et al., 2014; Kaur et al., 2020; Gnedin & Madau, 2022). For instance, Becker et al. (2011) reported the discovery of a very long and very opaque Gunn-Peterson trough in an otherwise mostly ionized Universe. Extreme features like this one can not be captured in our simulations (see e.g. Keating et al. 2020 for an estimation of the volume needed to simulate similar features). Therefore, it is possible that our results will be marginally affected by a somewhat suppressed variability due to the inability of thesan to capture the most extreme features. However, we expect these to be very rare, and therefore to bear a small impact on our results.
@@ 2.2 Synthetic spectra
The main data product used in this paper are synthetic lines of sights (LOS) extracted from the simulation outputs (see Section 3.10 of Garaldi et al., 2024). These are extracted using the colt code (last described in Smith et al., 2022b), which uses the native Voronoi tessellation of the simulation to retain the full spatial information available. Using the gas properties extracted in this way, we construct synthetic Ly forest spectra using the full Voigt-Hjerting line profile (Hjerting 1938, through the approximation of Harris 1948 and Tepper-García 2006). We include the effects of gas temperature and peculiar velocities. Our spectra have a spectral resolution of . For each simulation snapshot (i.e. approximately every 11 Myr) having redshift between , we produce 300 lines of sight (in addition to the 150 already available in the public data release). These new LOS originate from a random location in the plane, are aligned along the direction and are long. We note here that, thanks to the unstructured mesh of arepo, the synthetic spectra do not suffer from any grid-alignment artefacts.
For this paper, we combine the LOS produced at each simulation snapshot to produce lightcone spectra, employing a piecewise-constant approximation. We release these sightlines on the thesan website (see Appendix A for more information). In Fig. 1 we show an example of such lightcone LOS. The background map shows the density distribution within 30 Mpc of the sightline (along the direction of the simulation, left-hand side vertical axis and top left colorbar). The saturation of this map reflects the amount of neutral hydrogen in the IGM (with grey regions corresponding to and progressively more saturated colors indicating lower neutral fractions, top right colorbar). The white spectrum super-imposed to this map shows the normalised transmitted flux along the LOS (right-hand side y axis). The evolving redshift of the spectrum is indicated by the horizontal axis. This visually shows the well-known boost in Ly transmission in highly-ionized gas, which is preferentially found at low redshift.
2.3 Computing the GaLCC
Here, we briefly summarize how the GaLCC is computed in the paper. Given a set of spectra and a set of galaxy positions, we first compute the distance between each pixel in each spectrum and each galaxy position . Then, we discretize these distances into bins and compute the average transmitted flux for all pixel-galaxy pairs within each bin. Finally, we divide the result by the average transmitted flux of the entire sample of spectra (i.e. using every pixel in every spectrum) and subtract one to adhere to the convention in the literature. Notice that, by virtue of the fact that the synthetic spectra are randomly sampling the entire volume and are many in number, this procedure is equivalent to normalizing by the average Ly transmitted flux in the Universe, as typically done in observations where the number of available spectra is low and the procedure described might yield biased results because of sample variance.
In the paper, when not stated otherwise, the cross-correlation is always computed at using all the 300 sightlines (either at fixed redshift or using the lightcone LOS) and considering galaxies with stellar mass (corresponding to 1834 galaxies in the simulation box at ). We use 100 bins in the radial direction, linearly-spaced between and . Additionally, all quantities are expressed in comoving units, unless specified otherwise.